OpenStudy (anonymous):

(4b-2)(5b-4) would this = 9b-6

OpenStudy (preetha):

Multiplying, 20b^2-26b+8. Looks like a quadratic to me.

OpenStudy (anonymous):

Use Distributive Law

OpenStudy (anonymous):

Use the foil method. Good luck,

OpenStudy (anonymous):

(a+b) * (d-c) = ad - ac + bd-bc

OpenStudy (anonymous):

I did you the FOIL method and that is how I got 9b-6... is that wrong?

OpenStudy (anonymous):

\[(4b-2)(5b-4) = 20b^2 -16b - 10b + 8\]

OpenStudy (anonymous):

\[=20b^2-26b+8\]

OpenStudy (anonymous):

Like Preetha Said it would be a Quad Eq

OpenStudy (preetha):

Yahoo, you have a good Math background!

OpenStudy (anonymous):

okay I re-did it.. and I see what I did wrong... thank you Preetha and Yahoo for the help!

OpenStudy (anonymous):

Preetha - are you still there?

OpenStudy (preetha):

Yes